# Yuhang ${}^{\spadesuit}$   Siyuan ${}^{\clubsuit}$   Shujun ${}^{\clubsuit}$ Xuanjing ${}^{\diamondsuit}$   Zhongyu ${}^{\clubsuit\triangle}$

### 11 Mar 2024 v1

@@ Abstract
We introduce ALaRM, the first framework modeling hierarchical rewards in reinforcement learning from human feedback (RLHF), which is designed to enhance the alignment of large language models (LLMs) with human preferences. The framework addresses the limitations of current alignment approaches, which often struggle with the inconsistency and sparsity of human supervision signals, by integrating holistic rewards with aspect-specific rewards. This integration enables more precise and consistent guidance of language models towards desired outcomes, particularly in complex and open text generation tasks. By employing a methodology that filters and combines multiple rewards based on their consistency, the framework provides a reliable mechanism for improving model alignment. We validate our approach through applications in long-form question answering and machine translation tasks, employing gpt-3.5-turbo for pairwise comparisons, and demonstrate improvements over existing baselines. Our work underscores the effectiveness of hierarchical rewards modeling in refining LLM training processes for better human preference alignment. We release our codes at https://ALaRM-fdu.github.io.
@@
Yuhang ${}^{\spadesuit}$   Siyuan ${}^{\clubsuit}$   Shujun ${}^{\clubsuit}$ Xuanjing ${}^{\diamondsuit}$   Zhongyu ${}^{\clubsuit\triangle}$$^{\spadesuit}$nstitute of Science and Technology for Brain-inspired Intelligence, Fudan University ${}^{\clubsuit}$chool of Data Science, Fudan University ${}^{\diamondsuit}$chool of Computer Science, Fudan University ${}^{\triangle}$esearch Institute of Intelligent and Complex Systems, Fudan University yhlai23@m.fudan.edu.cn, {wangsy18,xjhuang,zywei}@fudan.edu.cn
@@ 6 Related Work
Superhuman models should be capable of handling complex and creative tasks beyond human expertise (Burns et al., 2023), which raises the increasingly important issue of scalable oversight: how to provide reliable supervision signals within limited human capabilities (Amodei et al., 2016; Bowman et al., 2022)? Current methods include improving evaluation quality through human-AI collaboration (Irving et al., 2018; Christiano et al., 2018) and simplifying tasks into sub-tasks for more reliable assessments (Zhong et al., 2023; Lightman et al., 2023; Wu et al., 2021). Our framework adopts the latter approach, identifying simpler aspects for evaluation.

### 16 Mar 2024 v2

@@ Abstract
We introduce ALaRM, the first framework modeling hierarchical rewards in reinforcement learning from human feedback (RLHF), which is designed to enhance the alignment of large language models (LLMs) with human preferences. The framework addresses the limitations of current alignment approaches, which often struggle with the inconsistency and sparsity of human supervision signals, by integrating holistic rewards with aspect-specific rewards. This integration enables more precise and consistent guidance of language models towards desired outcomes, particularly in complex and open text generation tasks. By employing a methodology that filters and combines multiple rewards based on their consistency, the framework provides a reliable mechanism for improving model alignment. We validate our approach through applications in long-form question answering and machine translation tasks, employing gpt-3.5-turbo for pairwise comparisons, and demonstrate improvements over existing baselines. Our work underscores the effectiveness of hierarchical rewards modeling in refining LLM training processes for better human preference alignment. We release our code at https://ALaRM-fdu.github.io.
@@
Yuhang ${}^{\spadesuit}$   Siyuan ${}^{\clubsuit}$   Shujun ${}^{\clubsuit}$ Xuanjing ${}^{\diamondsuit}$   Zhongyu ${}^{\clubsuit\triangle}$ [^2] Corresponding author ${}^{\spadesuit}$nstitute of Science and Technology for Brain-inspired Intelligence, Fudan University ${}^{\clubsuit}$chool of Data Science, Fudan University ${}^{\diamondsuit}$chool of Computer Science, Fudan University ${}^{\triangle}$esearch Institute of Intelligent and Complex Systems, Fudan University yhlai23@m.fudan.edu.cn, {wangsy18,xjhuang,zywei}@fudan.edu.cn
@@ 6 Related Work
Superhuman models should be capable of handling complex and creative tasks beyond human expertise (Burns et al., 2023), which raises the increasingly important issue of scalable oversight: how to provide reliable supervision signals within limited human capabilities (Amodei et al., 2016; Bowman et al., 2022)? Current methods include improving evaluation quality through human-AI collaboration (Irving et al., 2018; Christiano et al., 2018) and simplifying tasks into sub-tasks for more reliable assessments (Wu et al., 2021; Zhong et al., 2023; Lightman et al., 2023; Liu and Alahi, 2024). Our framework adopts the latter one, identifying simpler aspects for evaluation.